OpenStudy (anonymous):

Evaluate the following geometric sum. 1/3 + 1/6 + 1/12 + 1/24 + ..... + 1/192

OpenStudy (anonymous):

i messed that up - method was right but i miscalculated

OpenStudy (anonymous):

the correct answer in 127/192

OpenStudy (anonymous):

sum = (1/3) * 1 - r^n ------ 1 - r = (1/3) * 1 - (1/2)^7 --------- 1 - r = (1/3) * (1 - 1/128) ---------- (1/2) = 127 * 2 127 ------ = --- = 127/192 128 * 3 64 * 3

OpenStudy (anonymous):

i did the same thing but i kept getting the wrong answer i think it's cause i put in my calculator the wrong way

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

no probs